Does lovastatin (a cholesterol-lowering drug) reduce the risk of heart attack? In a Texas study,researchers gave lovastatin to 2,325 people and an inactive substitute to 2,081 people (average age 58). After 5 years, 57 of the lovastatin group had suffered a heart attack, compared with 97 for the
inactive pill. (a) State the appropriate hypotheses. (b) Obtain a test statistic and p-value. Interpret the results at α = .01. (c) Is normality assured? (d) Is the difference large enough to be important?
(e) What else would medical researchers need to know before prescribing this drug widely?

Answer by stanbon(75887)

Does lovastatin (a cholesterol-lowering drug) reduce the risk of heart attack? In a Texas study,researchers gave lovastatin to 2,325 people and an inactive substitute to 2,081 people (average age 58).
After 5 years, 57 of the lovastatin group had suffered a heart attack, compared with 97 for the inactive pill.
-------------------------------------
(a) State the appropriate hypotheses.
Ho: p(placebo)-p(active) =0
Ha: p(placebo)-p(active) >0
--------------------------------
57/2325 = 0.025 97/2081 = 0.047
(b) Obtain a test statistic and p-value.
z(0.047-0.025) = (0.022)/sqrt[(0.047*0.953)/2081 + (0.025*0.975)/2325)]=3.889
---
p-value = P(z>3.889) = 5.043x10^-5
Interpret: There is a 0.00000005043% probability that test results could
have provided stronger evidence for rejecting Ho.
-----------------------------------------------------------
Interpret the results at α = .01.
Since p-value is less than 1% reject Ho.
The lovastatin is a better medicine than the placebo.
-----------------------------------------------------
(c) Is normality assured?
Check whatever normality conditions are required by your text.
------------------------------------------------------
(d) Is the difference large enough to be important?
It is enormous.
-------------------------------
(e) What else would medical researchers need to know before prescribing this drug widely?
I'll leave that to you.
======================================
Cheers,
Stan H.